OpenStudy (anonymous):

how would i write (8,1),(2,3) in a equation midpoint what number would i substitute

OpenStudy (anonymous):

x+x/2 , y+y/2 8+2=10/2=5 1-3=-2/2=-1 (5,-1)

OpenStudy (anonymous):

the answer is (5,2) but when itry to do substitute part im confused

OpenStudy (anonymous):

y=mx+b find slope with the two points y-y/x-x input slope for m input x and y for their places in the y=mx+b equation and solve for b then rewrite without the x and y. only inputing the m and b that you found

OpenStudy (anonymous):

thanks